OpenStudy (aonz):

Halp please. If x=log2 and y=log3 express the following in terms of x and y.... log5.4

OpenStudy (aonz):

@daniel987600 @RadEn help pls

OpenStudy (raden):

log 5.4 = log (2 * 27/10) = log 2 + log 27 - log 10 = log 2 + log 3^3 - log 10 = log 2 + 3log 3 - 1 = ...

OpenStudy (aonz):

ohhh i get it now :) You divide by ten to move decimal :) answer is 3y + x -1